Three charges of +5c , -10c and -8c are distributed as shown below.....
Calculate the resultant force acting on the 5c
1
Expert's answer
2020-12-04T07:32:37-0500

We have three charges: +5, -10, -8.

the formula of force acting between two charges is F = "\frac{ k*q1*q2}{r2}" ;


There are two charges whch are force acting with +5c. these are -10, -8. Let's consider the location of charges in series and the distance of two charges which are in series r.

F+5 = F1 + F2 ; k = 8.99x109 Nm2/C2.


F1= "\frac{8.99 * 5c * 10c}{ r2}" x109 Nm2/C2 = "\frac{4.495}{ r2}" * 1011  N.

F2= "\frac{8.99 * 5c * 8c}{ r2}" x109 Nm2/C2 = "\frac{3.596}{ r2}" * 1011 N.


F= F1 + F2 = "\frac{4.495}{ r2}" * 1011  N + "\frac{3.596}{ r2}" * 1011 N = "\frac{8.091}{ r2}" * 1011 N.